# A positive period derivative in the quasi-periodic eruptions of ZTF19acnskyy

### v1

@@ IV.4 Hierarchical SMBH binary
$\dot{P}_{\rm max}\approx 4.4\times 10^{-5}\;\bigg(\frac{M_{\rm tot}}{2\times 10^{6}M_{\odot}}\bigg)^{1/3}\bigg(\frac{P_{\rm SMBH}}{20\;\mathrm{yr}}\bigg)^{-4/3}\bigg(\frac{P}{11\;\mathrm{d}}\bigg)$ (16)
@@ V Conclusion
We reported the surprising result that the quasi-period eruptions (QPEs) in ZTF19acnskyy/“Ansky” have shown a smoothly increasing period throughout 2025-2026, consistent with a time-derivative $\dot{P}\approx(1.7\pm 0.02)\times 10^{-2}$ (Fig. 3). Backwards-extrapolating the $\dot{P}$ implies that the 4.574.5-7 day recurrence times observed in 2024 data are expected $\gtrsim 20$ QPEs before those in 2025, providing a partial explanation for the previous data. The period derivative has remained constant for over a full year thus far, ruling out period oscillations with $P\lesssim 11$ yr or amplitude $\lesssim 1000$ d. It remains possible that a very long-term sinusoidal modulation, of period $\sim 11-27$ yr and amplitude $\sim 1000-4500$ d, is responsible for the observed period increase, though that is about equally challenging to explain.
@@
    Prior work has calculated the orbital evolution due to pericenter mass-transfer in eccentric stellar-mass binaries via the orbit-averaged Lagrange planetary equations (e.g. Sepinsky et al. 2009; Dosopoulou and Kalogera 2016b). Extending these results to EMRIs, the observed $\dot{P}$ can be reproduced for fractional mass losses of $\Delta M_{*}\sim 10^{-3}M_{*}$ per passage (Eq. 5), even for non-conservative mass-transfer ($\dot{M}_{\rm tot},\dot{J}_{\rm orb}\neq 0$). Two uncertainties are whether the stellar structure can sustain many consecutive comparable-energy bursts after losing $\sim$10% of its mass over a year (Fig. 5), and whether the $\sim$58 day viscous time at pericenter can be reconciled with the $\lesssim 3$-day burst duration.
@@
    We also examined whether the period changes may be produced by general relativistic precession. We considered the effects of apsidal motion (Eqns. 13-15; Linial and Quataert 2024) and nodal precession of the orbiter/accretion disk (Franchini et al., 2023; Chakraborty et al., 2025a), finding the large $\dot{P}$ amplitude and the long duration over which it is stable are difficult to reproduce simultaneously (Figs. 7-8). Given that no known combination of relativistic precessions appears readily able to explain the data, models for precession-induced timing variations may require extension, a theme also seen in other recent QPE timing studies (Miniutti et al., 2025; Arcodia et al., in prep.).
@@ Acknowledgements
We thank Keith Gendreau, Zaven Arzoumanian, Scott Hughes, Eric Coughlin, Xin Pan, Daniel D’Orazio, Hengxiao Guo, and Zhen Yan for discussions and data which contributed to this work. We are particularly grateful to the NICER team for providing an exceptionally high-quality dataset which uniquely enabled this analysis. GM acknowledges support from grant n. PID2023-147338NB-C21 funded by Spanish MICIU/AEI/10.13039/501100011033 and ERDF/EU. JC acknowledges financial support from ANID – FONDECYT Regular 1251444, and Millennium Science Initiative Program NCN2023_0022023\_002. LHG acknowledges financial support from ANID program FONDECYT Iniciación 11241477. MG is funded by Spanish MICIU/AEI/10.13039/501100011033 and ERDF/EU grant PID2023-147338NB-C21.

### v2

@@ IV.4 Hierarchical SMBH binary
$\dot{P}\approx 4.4\times 10^{-5}\;\bigg(\frac{M_{\rm tot}}{2\times 10^{6}M_{\odot}}\bigg)^{1/3}\bigg(\frac{P_{\rm SMBH}}{20\;\mathrm{yr}}\bigg)^{-4/3}\bigg(\frac{P}{11\;\mathrm{d}}\bigg)$ (16)
@@ V Conclusion
We reported the surprising result that the quasi-period eruptions (QPEs) in ZTF19acnskyy/“Ansky” have shown a smoothly increasing period—exactly opposite to the gravitational wave and/or viscous drag-driven inspiral predicted by EMRI models—with a time-derivative $\dot{P}\approx(1.7\pm 0.02)\times 10^{-2}$ throughout 2025-2026 (Fig. 3). Backwards-extrapolating the $\dot{P}$ implies that the 4.574.5-7 day recurrence times observed in 2024 data are expected $\gtrsim 20$ QPEs before those in 2025, providing a partial explanation for the previous data. The period derivative has remained constant for over a full year thus far, ruling out period oscillations with $P\lesssim 11$ yr or amplitude $\lesssim 1000$ d. It remains possible that a very long-term sinusoidal modulation, of period $\sim 11-27$ yr and amplitude $\sim 1000-4500$ d, is responsible for the observed period increase, though that is about equally challenging to explain.
@@
    Prior work has calculated the orbital evolution due to pericenter mass-transfer in eccentric stellar-mass binaries via the orbit-averaged Lagrange planetary equations (e.g. Sepinsky et al. 2009; Dosopoulou and Kalogera 2016b). Extending these results to EMRIs, the observed $\dot{P}$ can be reproduced for fractional mass losses of $\Delta M_{*}\sim 10^{-3}M_{*}$ per passage (Eq. 5), even for non-conservative mass-transfer ($\dot{M}_{\rm tot},\dot{J}_{\rm orb}\neq 0$). Three uncertainties are whether the stellar structure can sustain many consecutive comparable-energy bursts after losing $\sim$10% of its mass over a year (Fig. 5); whether the $\sim$58 day viscous time at pericenter can be reconciled with the $\lesssim 3$-day burst duration; and the extent to which L2 mass loss modifies the expected period evolution due to mass-transfer.
@@
    We also examined whether the period changes may be produced by general relativistic precession. We considered the effects of apsidal motion (Eqns. 13-15; Linial and Quataert 2024) and nodal precession of the orbiter/accretion disk (Franchini et al., 2023; Chakraborty et al., 2025a), finding the large $\dot{P}$ amplitude and the long duration over which it is stable are difficult to reproduce simultaneously (Figs. 7-8). Given that no known combination of relativistic precessions appears readily able to explain the data, models for precession-induced timing variations may require extension, a theme also seen in other recent QPE timing studies (Miniutti et al., 2025; Arcodia et al., subm.).
@@ Acknowledgements
We thank Keith Gendreau, Zaven Arzoumanian, Scott Hughes, Eric Coughlin, Xin Pan, Daniel D’Orazio, Hengxiao Guo, and Zhen Yan for discussions and data which contributed to this work. We thank the anonymous referee for constructive comments which improved the manuscript. We are particularly grateful to the NICER team for providing an exceptionally high-quality dataset which uniquely enabled this analysis. GM acknowledges support from grant n. PID2023-147338NB-C21 funded by Spanish MICIU/AEI/10.13039/501100011033 and ERDF/EU. JCu acknowledges financial support from ANID – FONDECYT Regular 1251444, and Millennium Science Initiative Program NCN2023_0022023\_002. LHG acknowledges financial support from ANID program FONDECYT Iniciación 11241477. MG is funded by Spanish MICIU/AEI/10.13039/501100011033 and ERDF/EU grant PID2023-147338NB-C21. The version of QPE-FIT (v0.1.11) used in this work can be found at the static repository doi:10.5281/zenodo.18943053.